Calculate the standard deviation of the returns on Andrew’s Violins stock if projections include the following?

State of Economy Probability of State Economy Rate of Return if State Occurs
Boom 30% 15%
Normal 65% 12%
Recession 5% 6%
Answer #1

Expected return=Respective return*Respective probability

=(0.3*15)+(0.65*12)+(0.05*6)=12.6%

probability Return probability*(Return-Expected Return)^2
0.3 15 0.3*(15-12.6)^2=1.728
0.65 12 0.65*(12-12.6)^2=0.234
0.05 6 0.05*(6-12.6)^2=2.178
Total=4.14%

Standard deviation=[Total probability*(Return-Expected Return)^2/Total probability]^(1/2)

=(4.14)^(1/2)

=2.03%(Approx).